OpenStudy (anonymous):

which value is closest to the value of 23 √ ? a.3.46 b.-1 c.6 d.4.24

OpenStudy (anonymous):

d

OpenStudy (anonymous):

i assume u said closest to sqrt(23)

OpenStudy (dumbcow):

3^2 = 9 4^2 = 16 5^2 = 25 well 23 is between 16 and 25 therefore sqrt(23) is between 4 and 5 :)

OpenStudy (anonymous):

no i meant to say \[ 2\sqrt{3}\]

OpenStudy (dumbcow):

oh ok well kinda same thing 3 is between 1 and 4 so sqrt(3) is between 1 and 2 Multiply by 2 2sqrt(3) is between 2 and 4

OpenStudy (dumbcow):

im assuming you can't just use a calculator ??

OpenStudy (anonymous):

dont have one :(

OpenStudy (dumbcow):

graphcalc.com or windows has one for sure

OpenStudy (anonymous):

\[\sqrt{1}<\sqrt{3}<\sqrt{4}\] \[1<\sqrt{3}<2\] \[2<2\sqrt{3}<4\] Answer 3.46